Jackie has five red cars, three of which are Jaguars, and five black cars, one of which is a Jaguar. Her butler selects two of these cars at random.


What is the probability one is red and one is black?



What is the probability both are Jaguars, given that one is red and one is black?



What is the probability both are red, given that both are Jaguars?



What is the probability both are red, given that both are the same color?


Enter your answers as whole numbers or fractions in lowest terms.

Answer by Boreal(15235)

5 red cars, 3 jags
5 black cars, 1 jag
probability 1 red 1 black is (5/10)(5/9), (5/10)(5/9)=5/9
check by probability both red (5/10)(4/9)=2/9 plus probability both black, which is also 2/9 Probability NOT 1 red 1 black is 4/9 and 1-4/9=5/9
Given one is red and one is black,
(3/5)(1/5)=3/25
Given that both are jags (n=4), probability both are red is 3/4
Given both are the same color, probability they are red is 1/2 as shown above.